# Asymptotic Distribution-free Change-point Detection for Modern Data Based on a New Ranking Scheme

### v3

@@ Abstract
Change-point detection (CPD) involves identifying distributional changes in a sequence of independent observations. Among nonparametric methods, rank-based methods are attractive due to their robustness and effectiveness and have been extensively studied for univariate data. However, they are not well explored for high-dimensional or non-Euclidean data. This paper proposes a new method, Rank INduced by Graph Change-Point Detection (RING-CPD), which utilizes graph-induced ranks to handle high-dimensional and non-Euclidean data. The new method is asymptotically distribution-free under the null hypothesis, and an analytic $p$-value approximation is provided for easy type-I error control. Simulation studies show that RING-CPD effectively detects change points across a wide range of alternatives and is also robust to heavy-tailed distribution and outliers. The new method is illustrated by the detection of seizures in a functional connectivity network dataset, changes of digit images, and travel pattern changes in the New York City Taxi dataset.
Index Terms:
Graph-induced ranks; Tail probability; High-dimensional data; Network data.
@@ I Introduction
Given a sequence of independent observations, an important problem is to decide whether the observations are from the same distribution or there is a change of distribution at a certain time point. Change-point detection (CPD) has attracted a lot of interest since the seminal work of [1], evolving into two primary strands: online (or sequential) CPD, which aims to detect distributional changes in real-time as data flows in (see for example [2, 3, 4, 5, 6, 7]), and offline CPD, which involves analyzing a completely observed data sequence and is the focal point of this paper. In this big data era, CPD has diverse applications in many fields, including functional magnetic resonance recordings [8, 9], healthcare [10, 11], communication network evolution [12, 13, 14], and financial modeling [15, 16]. Parametric approaches (see for example, [17, 18, 19, 20, 21]) are useful to address the problem for univariate and low-dimensional data, however, they are limited for high-dimensional or non-Euclidean data due to a large number of parameters to be estimated unless strong assumptions are imposed.
@@
A few nonparametric methods have been proposed, including kernel-based methods [2, 22, 23, 24, 25], interpoint distance-based methods [26, 27] and graph-based methods [28, 29, 30, 31, 32, 33, 34]. When dealing with high-dimensional data, the curse of dimensionality can pose a huge problem. The kernel-based method [22] assumes that the high-dimensional data resides on a low-dimensional manifold to get around of this problem. On the other hand, the graph-based methods [30] incorporate a useful pattern caused by the curse of dimensionality [35], and are effective in analyzing high-dimensional and non-Euclidean data for a wide range of changes. However, the graph-based methods focused on unweighted graphs, which may cause information loss. Later, [27] adopted a similar idea and proposed an asymptotic distribution-free approach utilizing all interpoint distances that worked well for detecting both location and scale changes. However, their test statistics are time and memory-consuming and implicitly require the existence of the second moment of the underlying distribution, which can be violated by heavy-tailed data or outliers that are common in many applications.
@@
Among the nonparametric methods, rank-based methods are attractive for univariate data due to their robustness and effectiveness [36, 37, 38, 39, 40, 41, 42, 43]. However, they are less explored for high-dimensional or non-Euclidean data. Specifically, existing multivariate rank-based methods are limited in many ways. For instance, [44] proposed to use the component-wise rank, which requires the dimension of the data to be smaller than the number of observations and suffers from dependent covariates. [45] and [46] proposed the spatial rank-based methods, which were designed mainly for detecting mean shifts. [47] proposed to use the ranks obtained from data depths, which is often used for low-dimensional data and is computation-extensive when the dimension is high.
@@
Noticing the gap between the potential benefit of the rank-based method and the scarce exploration for multivariate/high-dimensional data, we propose a new rank-based method called Rank INduced by Graph Change-Point Detection (RING-CPD), which can be applied to high-dimensional and non-Euclidean data. Unlike previous works dealing with the ranks of observations that are often limited to low-dimensional distributions, we propose to use the rank induced by similarity graphs that can be applied to data whose dimension could be much larger than the sample size. The graph-induced rank [48] is the rank defined in the similarity graphs. Instead of treating all edges in the graph equally, we assign the rank as weights to each edge and construct the scan statistic based on the ranks. Discussions on this rank and the new test are presented in Section II. We prove that the proposed scan statistic is asymptotic distribution-free, facilitating its usage to a broader community of researchers and analysts. They are also consistent against all types of changes under certain conditions of the similarity graph (Section III). The proposed statistic can work for a wide range of alternatives and is robust to heavy-tailed distribution and outliers, as illustrated by extensive simulation studies in Section IV and three real data examples in Section V. The details of proofs of the theorems are deferred to the Supplementary Material.
@@ II Method
For a sequence of indeendent $\{y_{i}\}_{i=1}^{n}$, we consider testing
@@
$H_{2}:\exists 1\leq\tau_{1}<\tau_{2}\leq n,\quad y_{i}\sim\begin{cases}F_{1},&i=\tau_{1}+1,\ldots,\tau_{2},\\ F_{0},&\text{ otherwise,}\end{cases}$ (2)
where $F_{0}$ and $F_{1}$ are two different distribution. When there are multiple change points, the two alternatives can be applied recursively. Alternatively, our method in the following can be extended similarly to [49] to accommodate multiple change-points, using the idea of wild binary segmentation [50] or seeded binary segmentation [51].
Our methodology begins with transforming the observations $\{y_{i}\}_{i=1}^{n}$ into a graph-induced rank matrix $\mathbf{R}=(R_{ij})_{i,j=1}^{n}\in\mathbb{R}^{n\times n}$, encapsulating the similarity among observations. Then we derive scan statistics utilizing $\mathbf{R}$. We follow notations in [48]. For any two graphs $G_{1}$ and $G_{2}$ sharing identical vertex sets, we consider $G_{1}\cap G_{2}=\emptyset$ when there are no shared edges between them, and define $G_{1}\cup G_{2}$ as their unified graph with combined edges. Given $\{y_{i}\}_{i=1}^{n}$, we sequentially construct a series of simple similarity graphs [^1] A simple graph is a graph without self-loops and multiple edges between any two vertices. $\{G_{l}\}_{l=0}^{k}$, starting with an edgeless graph $G_{0}$. The process continues with:
$G_{l+1}=G_{l}\cup G_{l+1}^{*}\,,\text{ where }G_{l+1}^{*}=\arg\max_{G^{\prime}\in\mathcal{G}_{l+1}}\sum_{(i,j)\in G^{\prime}}S(Z_{i},Z_{j})\text{ and }\mathcal{G}_{l+1}=\{G^{\prime}\in\mathcal{G}:G^{\prime}\cap G_{l}=\emptyset\}\,.$
[…]
@@
where for an event $A$, $\mathbbm{1}(A)$ is an indicator function that equals to one if event $A$ occurs, and equals to zero otherwise. The graph-induced ranks impose more weights on the edges with higher similarity, thus incorporating more similarity information than the unweighted graph. In the meantime, the robustness property of the ranks makes the weights less sensitive to outliers compared to direct using distance. The behavior of the graph-induced rank matrix in the $10$-NNG is illustrated in Figure 2 by simulated data. The graph-induced rank depends implicitly on $k$, and we will discuss its choice in Section IV-A.
@@
The computational times for $k$-NNG and $k$-MST are $O(n^{2}d)$ and $O\big{(}n^{2}(d+\log n)\big{)}$ [52], respectively. For more computationally efficient graphs, the approximate $k$-NNG [53] can be used, whose computational complexity is $O\big{(}nd(\log n+k\log d)\big{)}$, which is usually faster than $O(n^{2}d)$ when $k$ is small.
After constructing the graph-included rank matrix $\mathbf{R}$, we are ready to propose the test statistics. For testing the changed interval alternative $H_{2}$ (2), each possible interval $(t_{1},t_{2}]$ for $1\leq t_{1}<t_{2}\leq n$ partitions the observations into two groups: one group containing all observations observed during $(t_{1},t_{2}]$, and the other group containing all observations observed outside of this interval. Then, for any candidate changed interval $(t_{1},t_{2}]$, we define two basic quantities based on the ranks:
$U_{1}(t_{1},t_{2})=\sum_{i=1}^{n}\sum_{j=1}^{n}R_{ij}\mathbbm{1}(t_{1}<i,j\leq t_{2})\text{ and }U_{2}(t_{1},t_{2})=\sum_{i=1}^{n}\sum_{j=1}^{n}R_{ij}\mathbbm{1}(i,j\leq t_{1}\text{ or }i,j>t_{2})\,.$ (3)
@@
$Z_{w}(t_{1},t_{2})=\frac{U_{w}(t_{1},t_{2})-{\mathbbm{E}}\big{(}U_{w}(t_{1},t_{2})\big{)}}{\sqrt{{\rm Var}\big{(}U_{w}(t_{1},t_{2})\big{)}}}\text{ and }Z_{\mathrm{diff}}(t_{1},t_{2})=\frac{U_{\mathrm{diff}}(t_{1},t_{2})-{\mathbbm{E}}\big{(}U_{\mathrm{diff}}(t_{1},t_{2})\big{)}}{\sqrt{{\rm Var}\big{(}U_{\mathrm{diff}}(t_{1},t_{2})\big{)}}}\,,$
@@
$U_{w}(t_{1},t_{2})=\frac{n-t_{2}+t_{1}-1}{n-2}U_{1}(t_{1},t_{2})+\frac{t_{2}-t_{1}-1}{n-2}U_{2}(t_{1},t_{2})\,.$
Here we use ${\mathbbm{E}}$, ${\rm Var}$, and ${\rm Cov}$ to respectively denote the expectation, variance, and covariance under the permutation null distribution, which places $1/n!$ probability on each of the $n!$ permutations of the order of the observations. Under the alternative hypothesis, it is possible that (i) both $U_{1}(t_{1},t_{2})$ and $U_{2}(t_{1},t_{2})$ are larger than their null expectations (a typical scenario under location alternatives) and (ii) one of them is larger than while the other one is smaller than its corresponding null expectation (a typical scenario under scale alternatives). See [35] for more discussions on these scenarios. For (i), $Z_{w}(t_{1},t_{2})$ will be large and for (ii), $|Z_{\mathrm{diff}}(t_{1},t_{2})|$ will be large. Thus, $M_{R}$ is powerful for different types of alternatives.
@@
Although the similarity measure $S$ inherently exhibits symmetry, the rank matrix $\mathbf{R}$ may not always maintain this property, particularly in scenarios where the graph is directed such as the $k$-NNG. For derivation simplicity, we symmetrize $\mathbf{R}$ by $0.5(\mathbf{R}+\mathbf{R}^{\scriptscriptstyle\sf T})$. Note that this symmetrization does not change the value of $U_{1}$ and $U_{2}$. Without further specialization, we use $\mathbf{R}$ to denote its symmetrized version in the following. The explicit expressions of ${\mathbbm{E}}\big{(}U_{1}(t_{1},t_{2})\big{)}$, ${\mathbbm{E}}\big{(}U_{2}(t_{1},t_{2})\big{)}$ and $\bm{\Sigma}(t_{1},t_{2})$ are presented in Lemma 1. Let
$\bar{R}_{i\cdot}=\frac{\sum_{j=1}^{n}R_{ij}}{n-1}\,,r_{0}=\frac{\sum_{i=1}^{n}\bar{R}_{i\cdot}}{n}\,,r_{1}^{2}=\frac{\sum_{i=1}^{n}\bar{R}_{i\cdot}^{2}}{n}\,,r_{d}^{2}=\frac{\sum_{i=1}^{n}\sum_{j=1}^{n}R_{ij}^{2}}{n(n-1)}\,,V_{d}=r_{d}^{2}-r_{0}^{2}\,,V_{r}=r_{1}^{2}-r_{0}^{2}\,.$
@@
$f_{1}(t)=\frac{2t(t-1)(n-t)(n-t-1)}{(n-2)(n-3)}\text{ and }f_{2}(t)=\frac{4t(n-t)(t-1)(t-2)(n-1)}{(n-2)(n-3)}\,.$
The proof of Lemma 1 is through combinatorial analysis. It can be done similarly to the proof of Theorem $2.1$ in [48] and thus omitted here. When testing against $H_{1}$, we reject $H_{0}$ if the scan statistic
@@
Change-point detection is closely related to two-sample hypothesis testing. Our proposed methods utilize two-sample test statistics $M_{R}$, $Z_{w}$, $Z_{\rm diff}$ introduced in a prior work [48]. However, change-point detection is significantly more complicated. For two-sample hypothesis testing, the group labels of the observations are provided. In contrast, for change-point detection, this information is unavailable due to the unknown change-point location. In fact, even the existence of a change-point or not is unknown. Here, we use the maximum of the scan statistic as the test statistic; that is, we compute the two-sample test statistic for each potential change-point location, and we reject the null hypothesis of no change point if any of these statistics exhibit sufficient evidence. Understanding the distribution of the scan statistic is considerably more complicated than its corresponding quantity under the two-sample testing framework. We examine the asymptotic distribution of the scan statistic in Section III.

### v4

@@ Abstract
Change-point detection (CPD) involves identifying distributional changes in a sequence of independent observations. Among nonparametric methods, rank-based methods are attractive due to their robustness and effectiveness, and have been extensively studied for univariate data. However, they are not well explored for high-dimensional or non-Euclidean data. This paper proposes a new method, Rank INduced by Graph Change-Point Detection (RING-CPD), which utilizes graph-induced ranks to handle high-dimensional and non-Euclidean data. The new method is asymptotically distribution-free under the null hypothesis, and an analytic $p$-value approximation is provided for easy type-I error control. Simulation studies show that RING-CPD effectively detects change points across a wide range of alternatives and is also robust to heavy-tailed distribution and outliers. The new method is illustrated by the detection of seizures in a functional connectivity network dataset, changes in digit images, and travel pattern changes in the New York City Taxi dataset.
@@ I Introduction
Given a sequence of independent observations, an important problem is to decide whether the observations are from the same distribution or there is a change of distribution at a certain time point. Change-point detection (CPD) has attracted a lot of interest since the seminal work of [page1954continuous], evolving into two primary strands: online (or sequential) CPD, which aims to detect distributional changes in real-time as data flows in (see for example [desobry2005online, tartakovsky2008asymptotically, xie2013sequential, chen2019sequential, xie2020sequential, chu2022sequential]), and offline CPD, which involves analyzing a completely observed data sequence and is the focal point of this paper. In this big data era, CPD has diverse applications in many fields, including functional magnetic resonance recordings [barnett2016change, zambon2019change], healthcare [staudacher2005new, malladi2013online], communication network evolution [kossinets2006empirical, eagle2009inferring, peel2015detecting], and financial modeling [bai1998estimating, talih2005structural]. Parametric approaches (see for example, [srivastava1986likelihood, zhang2010detecting, siegmund2011detecting, chen2012parametric, wang2018change]) are useful to address the problem for univariate and low-dimensional data, however, they are limited for high-dimensional or non-Euclidean data due to a large number of parameters to be estimated unless strong assumptions are imposed.
@@
A few nonparametric methods have been proposed, including kernel-based methods [desobry2005online, li2015m, garreau2018consistent, arlot2019kernel, chang2019kernel], interpoint distance-based methods [matteson2014nonparametric, li2020asymptotic] and graph-based methods [chen2015graph, shi2017consistent, chu2019asymptotic, chen2019change, song2020asymptotic, liu2020fast, nie2021weighted, chen2023graph, chu2025tightness]. When dealing with high-dimensional data, the curse of dimensionality can pose a huge problem. The kernel-based method [li2015m] assumes that the high-dimensional data resides on a low-dimensional manifold to get around this problem. On the other hand, the graph-based methods [chu2019asymptotic] incorporate a useful pattern caused by the curse of dimensionality [chen2017new], and are effective in analyzing high-dimensional and non-Euclidean data for a wide range of changes. However, the graph-based methods have focused on unweighted graphs, which may cause information loss. Later, [li2020asymptotic] adopted a similar idea and proposed an asymptotic distribution-free approach utilizing all interpoint distances that worked well for detecting both location and scale changes. However, their test statistics are time and memory-consuming and implicitly require the existence of the second moment of the underlying distribution, which can be violated by heavy-tailed data or outliers that are common in many applications.
@@
Among the nonparametric methods, rank-based methods are attractive for univariate data due to their robustness and effectiveness [bhattacharyya1968nonparametric, darkhovskh1976nonparametric, pettitt1979non, schechtman1982nonparametric, lombard1987rank, lombard1983asymptotic, gerstenberger2018robust, wang2020rank]. However, they are less explored for high-dimensional or non-Euclidean data. Specifically, existing multivariate rank-based methods are limited in many ways. For instance, [lung2015homogeneity] proposed to use the component-wise rank, which requires the dimension of the data to be smaller than the number of observations and suffers from dependent covariates. [zhang2020spatial] and [shu2022spatial] proposed the spatial rank-based methods, which were designed mainly for detecting mean shifts. [chenouri2020robust] proposed to use the ranks obtained from data depths, which is often used for low-dimensional data and is computationally intensive when the dimension is high.
@@
Noticing the gap between the potential benefit of the rank-based method and the scarce exploration for multivariate/high-dimensional data, we propose a new rank-based method called Rank INduced by Graph Change-Point Detection (RING-CPD), which can be applied to high-dimensional and non-Euclidean data. Unlike previous works dealing with the ranks of observations that are often limited to low-dimensional distributions, we propose to use the rank induced by similarity graphs that can be applied to data whose dimension could be much larger than the sample size. The graph-induced rank [rise] is the rank defined in the similarity graphs. Instead of treating all edges in the graph equally, we assign the rank as weights to each edge and construct the scan statistic based on the ranks. Discussions on this rank and the new test are presented in Section II. We prove that the proposed scan statistic is asymptotically distribution-free, facilitating its usage to a broader community of researchers and analysts. They are also consistent against all types of changes under certain conditions of the similarity graph (Section III). The proposed statistic can work for a wide range of alternatives and is robust to heavy-tailed distributions and outliers, as illustrated by extensive simulation studies in Section IV and three real data examples in Section V. The details of proofs of theorems are deferred to the Supplementary Material.
@@ II Method
For a sequence of independent $\{y_{i}\}_{i=1}^{n}$, we consider testing
@@
$H_{2}:\exists 1\leq\tau_{1}<\tau_{2}\leq n,y_{i}\sim\begin{cases}F_{1},&i\in[\tau_{1}+1,\tau_{2}]\\ F_{0},&\text{otherwise,}\end{cases}$ (2)
where $F_{0}$ and $F_{1}$ are two different distributions. When there are multiple change points, the two alternatives can be applied recursively. Alternatively, our method in the following can be extended similarly to [zhang2021graph] to accommodate multiple change-points, using the idea of wild binary segmentation [fryzlewicz2014wild] or seeded binary segmentation [kovacs2020seeded].
Our methodology begins by transforming the observations $\{y_{i}\}_{i=1}^{n}$ into a graph-induced rank matrix $\mathbf{R}=(R_{ij})_{i,j=1}^{n}\in\mathbb{R}^{n\times n}$, which encapsulates the similarity between pairs of observations. Then we derive scan statistics utilizing $\mathbf{R}$. We follow notations in [rise]. Let $V=\{1,2,\ldots,n\}$ denote the vertex set of the $n$ observations. A graph $G$ on these observations is defined as $G=(V,E)$, where $E\subseteq V\times V$. We define that an edge $(i,j)\in G$ if $(i,j)\in E$. For two graphs, $G_{1}=(V,E_{1})$ and $G_{2}=(V,E_{2})$, sharing the same vertex set $V$, we define their intersection as $G_{1}\cap G_{2}=(V,E_{1}\cap E_{2})$, which is empty when there are no shared edges, i.e., $G_{1}\cap G_{2}=\emptyset$ if $E_{1}\cap E_{2}=\emptyset$, and their union as $G_{1}\cup G_{2}=(V,E_{1}\cup E_{2})$. Given $\{y_{i}\}_{i=1}^{n}$, we sequentially construct a series of simple similarity graphs [^1] A simple graph is a graph without self-loops and multiple edges between any two vertices. $\{G_{l}\}_{l=0}^{k}$, starting with an edgeless graph $G_{0}=(V,\emptyset)$. The process continues with $G_{l+1}=G_{l}\cup G_{l+1}^{*},$ where $G_{l+1}^{*}=\arg\max_{G^{\prime}\in\mathcal{G}_{l+1}}\sum_{(i,j)\in G^{\prime}}S(y_{i},y_{j})\text{ and }\mathcal{G}_{l+1}=\{G^{\prime}\in\mathcal{G}:G^{\prime}\cap G_{l}=\emptyset\}.$ Here, $\mathcal{G}$ represents a set of graphs subject to certain structural constraints, and $S(\cdot,\cdot)$ is a similarity measure, such as $S(y_{i},y_{j})=-\|y_{i}-y_{j}\|$ for Euclidean data, with $\|\cdot\|$ denoting the Euclidean norm. This framework allows for the construction of various well-established similarity graphs under different constraints. For instance, both the $k$-nearest neighbor graph ($k$-NNG) and the $k$-minimum spanning tree ($k$-MST) [^2] The MST is a spanning tree that minimizes the sum of distances of edges in the tree while connecting all observations. The $k$-MST is the union of the 11st, …, $k$th MSTs, where the $k$th MST is a spanning tree that connects all observations while minimizing the sum of distances across edges excluding edges in the (k1)(k-1)-MST.[friedman1979multivariate] satisfy the above definition of the sequence of graphs. For NNG, the structural constraint on $G^{\prime}$ is that each vertex $i$ connects to exactly one other vertex. Specifically, the graph set $\mathcal{G}$ is defined as $\mathcal{G}=\{G^{\prime}=(V,E^{\prime}):$ for each $i$, there exists one and only one $j\in V\backslash\{i\}$ such that $(i,j)\in E^{\prime}\}$. Thus, $G_{1}$ corresponds to the $1$-NNG, where each vertex $i$ connects to the vertex $j$ with the largest similarity $S(y_{i},y_{j})$ among all other observations. Similarly, $G_{l}$ is the $l$-NNG, $G_{l+1}^{*}$ is the $(l+1)$th NNG and $G_{l+1}$ is the $(l+1)$-NNG. For $k$-MST, the constraint is that $G^{\prime}$ must be a tree connecting all vertices, making $G_{l}$ the $l$-MST, $G_{l+1}^{*}$ the $(l+1)$th MST and $G_{l+1}$ the $(l+1)$-MST. An illustration of these graphs is presented in Figure 1. For more choices of graphs, one can see [rise].
@@
where for an event $A$, $\mathbbm{1}(A)$ is an indicator function that equals one if event $A$ occurs, and equals zero otherwise. The graph-induced ranks impose more weights on the edges with higher similarity, thus incorporating more similarity information than the unweighted graph. In the meantime, the robustness property of the ranks makes the weights less sensitive to outliers compared to directly using distance. The behavior of the graph-induced rank matrix in the $10$-NNG is illustrated in Figure 2 by simulated data. The graph-induced rank depends implicitly on $k$, and we will discuss its choice in Section IV-A.
@@
The computational complexities for constructing $k$-NNG and $k$-MST using brute-force methods are $O(n^{2}d)$ and $O\big(n^{2}(d+\log n)\big)$ [friedman1979multivariate], respectively. More efficient algorithms are widely used in practice. For example, the $k$-NNG can be computed using KD Tree [bentley1975multidimensional] or Ball Tree algorithms [omohundro1989five], as implemented in scikit-learn [JMLR:v12:pedregosa11a], with a computational complexity of $O(n\log n)$ when the dimension $d$ is small (e.g., $d\leq 15$ as recommended by scikit-learn package). However, these methods face challenges due to the “curse of dimensionality.” In such cases, both their empirical and theoretical performance can degrade, and may become comparable or even worse than brute-force algorithms, as discussed in [ram2019revisiting]. On the other hand, the approximate $k$-NNG offers a computational complexity of $O\big(nd(\log n+k\log d)\big)$, even with high-dimensional data [arya1998optimal, ram2019revisiting].
After constructing the graph-induced rank matrix $\mathbf{R}$, we are ready to propose the test statistics. For testing the changed interval alternative $H_{2}$ (2), each possible interval $(t_{1},t_{2}]$ for $1\leq t_{1}<t_{2}\leq n$ partitions the observations into two groups: one group containing all observations observed during $(t_{1},t_{2}]$, and the other group containing all observations observed outside of this interval. Then, for any candidate changed interval $(t_{1},t_{2}]$, we define two basic quantities:
$U_{1}(t_{1},t_{2})=\sum_{i=1}^{n}\sum_{j=1}^{n}R_{ij}\mathbbm{1}(t_{1}<i,j\leq t_{2}),$ (3)
$U_{2}(t_{1},t_{2})=\sum_{i=1}^{n}\sum_{j=1}^{n}R_{ij}\mathbbm{1}(i,j\leq t_{1}\text{ or }i,j>t_{2}).$ (4)
@@
$Z_{w}(t_{1},t_{2})=\frac{U_{w}(t_{1},t_{2})-{\mathbbm{E}}\big(U_{w}(t_{1},t_{2})\big)}{\sqrt{{\rm Var}\big(U_{w}(t_{1},t_{2})\big)}},$
$Z_{\mathrm{diff}}(t_{1},t_{2})=\frac{U_{\mathrm{diff}}(t_{1},t_{2})-{\mathbbm{E}}\big(U_{\mathrm{diff}}(t_{1},t_{2})\big)}{\sqrt{{\rm Var}\big(U_{\mathrm{diff}}(t_{1},t_{2})\big)}}\,$,
@@
$U_{w}(t_{1},t_{2})=\tfrac{n-t_{2}+t_{1}-1}{n-2}U_{1}(t_{1},t_{2})+\tfrac{t_{2}-t_{1}-1}{n-2}U_{2}(t_{1},t_{2}).$
Here we use $\mathbbm{E}$}, $\rm Var$}, and $\rm Cov$} to respectively denote the expectation, variance, and covariance under the permutation null distribution, which places $1/n!$ probability on each of the $n!$ permutations of the order of the observations. Under the alternative hypothesis, it is possible that (i) both $U_{1}(t_{1},t_{2})$ and $U_{2}(t_{1},t_{2})$ are larger than their null expectations (a typical scenario under location alternatives) and (ii) one of them is larger than while the other one is smaller than its corresponding null expectation (a typical scenario under scale alternatives). See [chen2017new] for more discussions on these scenarios. For (i), $Z_{w}(t_{1},t_{2})$ will be large and for (ii), $|Z_{\mathrm{diff}}(t_{1},t_{2})|$ will be large. Thus, $M_{R}$ is powerful for different types of alternatives.
@@
Although the similarity measure $S$ inherently exhibits symmetry, the rank matrix $\mathbf{R}$ may not always maintain this property, particularly in scenarios where the graph is directed such as the $k$-NNG. For simplicity, we symmetrize $\mathbf{R}$ by $0.5(\mathbf{R}+\mathbf{R}^{\scriptscriptstyle\sf T})$. Note that this symmetrization does not change the value of $U_{1}$ and $U_{2}$. Without further specialization, we use $\mathbf{R}$ to denote its symmetrized version in the following. The explicit expressions of ${\mathbbm{E}}\big(U_{1}(t_{1},t_{2})\big)$, ${\mathbbm{E}}\big(U_{2}(t_{1},t_{2})\big)$ and $\bm{\Sigma}(t_{1},t_{2})$ are presented in Lemma 1. Let
$\bar{R}_{i\cdot}=\frac{\sum_{j=1}^{n}R_{ij}}{n-1}\,,r_{0}=\frac{\sum_{i=1}^{n}\bar{R}_{i\cdot}}{n}\,,r_{1}^{2}=\frac{\sum_{i=1}^{n}\bar{R}_{i\cdot}^{2}}{n}\,$,
$r_{d}^{2}=\frac{\sum_{i=1}^{n}\sum_{j=1}^{n}R_{ij}^{2}}{n(n-1)}\,,V_{d}=r_{d}^{2}-r_{0}^{2}\,,V_{r}=r_{1}^{2}-r_{0}^{2}\,.$
@@
$f_{1}(t)=\frac{2t(t-1)(n-t)(n-t-1)}{(n-2)(n-3)},$
$f_{2}(t)=\frac{4t(n-t)(t-1)(t-2)(n-1)}{(n-2)(n-3)}.$
The proof of Lemma 1 is through combinatorial analysis. It can be done similarly to the proof of Theorem $2.1$ in [rise] and thus omitted here. When testing against $H_{1}$, we reject $H_{0}$ if the scan statistic
@@
Change-point detection is closely related to two-sample hypothesis testing. Our proposed methods utilize two-sample test statistics $M_{R}$, $Z_{w}$, $Z_{\rm diff}$ introduced in a prior work [rise]. However, change-point detection is significantly more complicated. For two-sample hypothesis testing, the group labels of the observations are provided. In contrast, for change-point detection, this information is unavailable due to the unknown change-point location. In fact, even the existence of a change-point or not is unknown. Here, we use the maximum of the scan statistic as the test statistic; that is, we compute the two-sample test statistic for each potential change-point location, and we reject the null hypothesis of no change point if any of these statistics exhibit sufficient evidence. Understanding the distribution of the scan statistic is considerably more complicated than its corresponding quantity under the two-sample testing framework. We examine the asymptotic distribution of the scan statistic in Section III.